# Kalman–Bucy Filtering with Randomized Sensing:

### v1

@@ I Introduction
Steady-state analysis of the estimation error reveals how uncertainty evolves and stabilizes over time [1]. In traditional settings, the Kalman filter (KF) provides the optimal linear minimum-variance estimator, with its steady-state covariance governed by a Riccati equation [2, 3, 4]. However, in modern applications such as estimating a spatiotemporal field using a multi-agent sensor network, the sensing process itself is inherently stochastic—sensor locations, measurement timing, and noise may all vary randomly [5, 6, 7]. This randomness makes direct steady-state analysis of the estimation error intractable. Consequently, it is more insightful to adopt a statistical viewpoint, by modeling sensing events as random processes and analyzing their collective impact on estimation performance. Within this stochastic sensing framework, we address a fundamental sensor network design problem constrained by performance requirements: How many sensors, and of what quality, are required to guarantee a prescribed level of estimation accuracy prior to deployment? This motivates two key objectives: (1) to characterize the expected estimation performance analytically, and (2) to establish the fundamental limits of estimation accuracy under stochastic sensing. In this work, we pursue these objectives through both theoretical analysis and application-driven validation.
@@
Prior work has addressed this question in the discrete-time setting. In particular, Gupta [8] analyzed the stochastic sensor selection problem and derived upper bounds on the expected error covariance of the KF under randomly varying measurement availability. While insightful, these bounds are recursive in nature and therefore computationally demanding. In this work, we approximate the measurement process using a continuous-time model rather than directly analyzing the discrete-time recursion. This formulation leads to a differential inequality governing the evolution of the expected estimation covariance, from which the asymptotic behavior of the estimation system naturally follows.
@@
The continuous-time approximation offers two main advantages. First, it admits a closed-form expression for the steady-state covariance, making the analysis computationally much more efficient than the recursive discrete-time formulation. Second, in the context of multi-agent estimation of spatiotemporal Gaussian Process (GP) fields [9], where agents collect noisy measurements at locations drawn from a spatial distribution, the GP structure maps naturally to a linear state-space model, allowing the KF to produce optimal estimates. To quantify estimation uncertainty within this framework, we adopt the clarity, a rescaled form of the differential entropy of a random variable [10]. Leveraging the derived closed-form upper bound and its asymptotic behavior, the proposed approach reveals that the lower bound of the averaged expected clarity is independent of the discretization grid. This property enables the identification of the fundamental limits that govern steady-state estimation performance under stochastic sensing. Consequently, the analysis provides a systematic tool for determining, prior to deployment, (i) the minimal number of agents and (ii) the required sensing capabilities to achieve a prescribed level of estimation accuracy. This directly addresses an optimal sensor resource design problem that cannot be solved efficiently within the discrete-time formulation.
@@ II Literature Review
A substantial body of work has analyzed the covariance of the KF estimation error, with particular emphasis on its steady-state behavior under various system conditions. For classical time-invariant systems, it is well established that under detectability and stabilizability, the covariance of estimation error converges to a steady state in both discrete- and continuous-time settings [11]. Ni et al. [12] further showed that, for systems with time-varying output matrices, uniform complete observability alone ensures asymptotic stability of the continuous-time KF. However, the systems described above assume deterministic sensing. In contrast, modern sensing environments often deviate from this assumption due to inherent randomness: packet losses, stochastic sensor scheduling, or random sensor placement [7, 8, 13].
@@
Sinopoli et al. [7] pioneered the study of the KF under random measurements by modeling packet losses as a Bernoulli process. They established statistical convergence properties of the expected covariance of estimation error and identified a critical observation-arrival rate, below which the covariance diverges and above which it remains bounded.
@@
Gupta et al. [8] extended this random measurement framework to a more general sensor scheduling problem. They proposed a stochastic sensor selection strategy that activates one sensor at each time step according to a prescribed probability distribution. Using the discrete-time Riccati recursion, they analyzed the evolution of the covariance of estimation error in expectation, deriving upper and lower bounds. Their results further demonstrated, through examples, that these bounds accurately approximate the actual steady-state covariance.
@@
Subsequent research generalized these results to the setting of random measurement matrices. For example, Marelli et al. [5] considered the case where both the measurement matrix and the noise covariance are drawn randomly at each time from a known distribution, and established necessary and sufficient conditions for mean-square stability of the KF. While closely related to our problem, their analysis is limited to stability criteria and does not provide bounds or steady-state characterizations of the error covariance.
@@
Calle and Bopardikar [13] derived probabilistic semidefinite bounds on the covariance of estimation error for randomized sensor selection in Kalman Filtering. Their framework, however, models sensor selection as a one-time random sampling of a subset of sensors that remains fixed thereafter, in contrast to the per-step stochastic scheduling analyzed by Gupta et al. [8].
@@
Early approaches primarily relied on the discrete-time Riccati recursion to track the distribution or expectation of the covariance of estimation error. More recently, efforts have been made to analyze continuous-time process models with discrete-time measurement models [14, 6, 15]. For example, Ahdab et al. [14] modeled measurement arrivals as independent Poisson processes with sensor-specific rates and derived an upper bound on the mean posterior covariance of the continuous–discrete KF along the mean auxiliary state. Similarly, Dasgupta and Tanwani [6] studied continuous-time linear systems with Poisson-sampled observations. They showed that the Kalman–Bucy filter yields piecewise-deterministic covariance trajectories, and derived conditions on the mean observation rate that ensure bounded and convergent expected covariance of estimation error. Their results extend the discrete-time “dropout threshold” in [7]. However, their setting assumes a constant measurement matrix, with randomness arising from the irregular Poisson-distributed observation times rather than from the measurement matrix itself. In contrast, our focus is on a more general problem, akin to the framework of Marelli et al. [5], where the measurement matrix itself is random.
@@
Related work has also addressed continuous measurement models. For example, [16] study sensor scheduling for linear time-invariant systems. The authors derive a convex relaxation of the scheduling problem, yielding a performance lower bound. A lower bound reflects only the best-case performance and offers no guarantee on how the actual performance compares. In contrast, an upper bound constrains the worst-case covariance of the estimation error, providing a concrete performance guarantee. Additionally, [17] consider informative planning of sensing agents over an infinite time horizon for continuous-time linear systems. The covariance bound derived in this work tends to be conservative, as it omits the contribution of the measurement term.
@@
Analysis of the covariance of the estimation error naturally informs the design of the sensing system. Once stability conditions or performance bounds have been established, a practical question arises: how should sensors be deployed or activated to guarantee a prescribed level of estimation accuracy? This leads to sensor scheduling formulations, in which the active subset is selected over time. Prior work has addressed these problems through design optimization grounded in covariance (or its bounds), deriving policies that guide per–time-step sensor activation [18, 14]. Another typical problem is optimal sensor placement: given system parameters, one may wish to determine the minimum number of sensors or the required sensing quality needed to achieve a prescribed uncertainty threshold [19]. Most prior work adopts covariance (or surrogates such as its log-determinant) as the performance metric and designs placement/scheduling policies accordingly. However, covariance measures are often difficult to interpret or compare across different sensing configurations, since they are unit-dependent and different scalarizations (e.g., trace, log-determinant) can rank designs inconsistently. To evaluate the estimation quality, we employ clarity, as introduced in [10, 9], which normalizes the differential entropy of a continuous random variable to the interval $[0,1]$. In this paper, we show that within our closed-form upper-bound framework, the fundamental limits of estimation performance can be explicitly characterized, enabling direct quantification of the sensing resources required to achieve a prescribed uncertainty threshold.
@@ III Problem Formulation
$\mathcal{R}_{\bm{G}}(\bm{\Sigma})=\bm{\bm{A}}\bm{\Sigma}+\bm{\Sigma}\bm{\bm{A}}^{\top}+\bm{Q}_{c}-\bm{\Sigma}\bm{G}\bm{\Sigma},$ (6)
@@ IV-A Continuous-Time Upper Bound of Expected Covariance
The operator $\mathcal{R}_{\bm{G}}$ is concave, that is, for any $\bm{\Sigma}_{1},\bm{\Sigma}_{2}\in\mathcal{S}_{++}^{n}$ and $\lambda\in[0,1]$,
$\mathcal{R}_{\bm{G}}(\lambda\bm{\Sigma}_{1}\!+\!(1-\!\lambda)\bm{\Sigma}_{2})\succeq\lambda\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{1})\!+\!(1-\!\lambda)\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{2})$ (9)
@@
$\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{1})-\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{2})=\tilde{\bm{A}}\bm{K}+\bm{K}\tilde{\bm{A}}^{\top}$ (10)
@@
$=\mathbb{E}[\mathcal{R}_{\bar{\bm{G}}}(\bm{\Sigma}(t))]$ (16d)
$\preceq\mathcal{R}_{\bar{\bm{G}}}(\mathbb{E}[\bm{\Sigma}(t)])$ (16e)
where Eq. (16a) holds by linearity of expectation; Eq. (16b) follows by substituting the expression for $\dot{\bm{\Sigma}}$; Eq. (16c) uses the assumed independence of $\bm{\Sigma}(t)$ and $\bm{G}(t)$; Eq. (16d) invokes the definition of the Riccati operator in Eq. (6), and Eq. (16e) follows from Jensen’s inequality together with the concavity of $\mathcal{R}_{\bm{G}}$ established in Lemma 1.
@@
$\succeq\mathcal{R}_{\bar{\bm{G}}}(\bm{\Delta}(t))-\mathcal{R}_{\bar{\bm{G}}}(\bar{\bm{\Sigma}}(t))$
@@ IV-B Closed-form Solution of Upper Bound
$0\;=\;a\bm{\Delta}_{\infty}+\bm{\Delta}_{\infty}a+q_{c}\bm{I}-\bm{\Delta}_{\infty}\bar{\bm{G}}\bm{\Delta}_{\infty},$ (22)
@@
Since $\bar{\bm{G}}\in\mathbb{S}_{++}^{n}$, it admits the eigen-decomposition $\bar{\bm{G}}=\bm{U}\Lambda\bm{U}^{\top}$, where $\bm{U}$ is orthogonal and $\Lambda$ is diagonal. Under the isotropic assumption $\bm{A}=a\bm{I}$ and $\bm{Q}_{c}=q_{c}\bm{I}$, Eq. (22) implies that $\bar{\bm{G}}$ and $\bm{\Delta}_{\infty}$ commute, i.e., $\bar{\bm{G}}\bm{\Delta}_{\infty}=\bm{\Delta}_{\infty}\bar{\bm{G}}$. By Theorem 4.5.15 (a) in [20], any two commuting symmetric matrices can be simultaneously diagonalized by the same orthogonal matrix. Hence, $\bm{\Delta}_{\infty}=\bm{U}\bm{\Gamma}\bm{U}^{\top}$, where $\bm{\Gamma}$ is diagonal.
Applying the same factorization to $\bm{\Delta}_{\infty}$, we have $\bm{\Delta}_{\infty}=\bm{U}\bm{\Gamma}\bm{U}^{\top}$, where $\bm{\Gamma}$ is diagonal. Therefore, the equation becomes
@@
$=2a(\bm{U}\bm{\Gamma}\bm{U}^{\top})+q_{c}\bm{I}-(\bm{U}\bm{\Gamma}\bm{U}^{\top})(\bm{U}\Lambda\bm{U}^{\top})(\bm{U}\bm{\Gamma}\bm{U}^{\top})$
$=q_{c}\bm{I}+\bm{U}(2a\bm{\Gamma}-\bm{\Gamma}\Lambda\bm{\Gamma})\bm{U}^{\top}$
$=\bm{U}(q_{c}\bm{I}+2a\bm{\Gamma}-\bm{\Gamma}\Lambda\bm{\Gamma})\bm{U}^{\top}$
where the term inside the parentheses is a diagonal matrix. For the equation to hold, each diagonal entry must be zero. Therefore, let $\gamma_{i}=\bm{\Gamma}_{ii}$, $\lambda_{i}=\Lambda_{ii}$, and we have
@@
Thus we have derived the eigenvalues of $\bm{\Delta}_{\infty}$. Using these eigenvalues, we have
$\bm{\Delta}_{\infty}=\bm{U}(\operatorname{diag}(\gamma_{1},...,\gamma_{n}))\bm{U}^{\top}$ (28)

### v2

@@ I Introduction
Steady-state analysis of the estimation error reveals how uncertainty evolves and stabilizes over time [anderson2005optimal]. In traditional settings, the Kalman filter (KF) provides the optimal linear minimum-variance estimator, with its steady-state covariance governed by a Riccati equation [grewal2025kalman, lan2016rapidly, seo2025kalman]. However, in modern applications such as estimating a spatiotemporal field using a multi-agent sensor network, the sensing process itself is inherently stochastic—sensor locations, measurement timing, and noise may all vary randomly [marelli2019stability, dasgupta2025steady, sinopoli2004kalman]. This randomness makes direct steady-state analysis of the estimation error intractable. Consequently, it is more insightful to adopt a statistical viewpoint, by modeling sensing events as random processes and analyzing their collective impact on estimation performance. Within this stochastic sensing framework, we address a fundamental sensor network design problem constrained by performance requirements: How many sensors, and of what quality, are required to guarantee a prescribed level of estimation accuracy prior to deployment? This motivates two key objectives: (1) to characterize the expected estimation performance analytically, and (2) to establish the fundamental limits of estimation accuracy under stochastic sensing. In this work, we pursue these objectives through both theoretical analysis and application-driven validation.
@@
Prior work has addressed this question in the discrete-time setting. In particular, Gupta [gupta2006stochastic] analyzed the stochastic sensor selection problem and derived upper bounds on the expected error covariance of the KF under randomly varying measurement availability. While insightful, these bounds are recursive in nature and therefore computationally demanding. In this work, we approximate the measurement process using a continuous-time model rather than directly analyzing the discrete-time recursion. This formulation leads to a differential inequality governing the evolution of the expected estimation covariance, from which the asymptotic behavior of the estimation system naturally follows.
@@
The continuous-time approximation offers two main advantages. First, it admits a closed-form expression for the steady-state covariance, making the analysis computationally much more efficient than the recursive discrete-time formulation. Second, in the context of multi-agent estimation of spatiotemporal Gaussian Process (GP) fields [agrawal2024multi], where agents collect noisy measurements at locations drawn from a spatial distribution, the GP structure maps naturally to a linear state-space model, allowing the KF to produce optimal estimates. To quantify estimation uncertainty within this framework, we adopt the clarity, a rescaled form of the differential entropy of a random variable [agrawal2023sensor]. Leveraging the derived closed-form upper bound and its asymptotic behavior, the proposed approach reveals that the lower bound of the averaged expected clarity is independent of the discretization grid. This property enables the identification of the fundamental limits that govern steady-state estimation performance under stochastic sensing. Consequently, the analysis provides a systematic tool for determining, prior to deployment, (i) the minimal number of agents and (ii) the required sensing capabilities to achieve a prescribed level of estimation accuracy. This directly addresses an optimal sensor resource design problem that cannot be solved efficiently within the discrete-time formulation.
@@ II Literature Review
A substantial body of work has analyzed the covariance of the KF estimation error, with particular emphasis on its steady-state behavior under various system conditions. For classical time-invariant systems, it is well established that under detectability and stabilizability, the covariance of estimation error converges to a steady state in both discrete- and continuous-time settings [lewis2017optimal]. Ni et al. [ni2016stability] further showed that, for systems with time-varying output matrices, uniform complete observability alone ensures asymptotic stability of the continuous-time KF. However, the systems described above assume deterministic sensing. In contrast, modern sensing environments often deviate from this assumption due to inherent randomness: packet losses, stochastic sensor scheduling, or random sensor placement [sinopoli2004kalman, gupta2006stochastic, calle2021probabilistic].
@@
Sinopoli et al. [sinopoli2004kalman] pioneered the study of the KF under random measurements by modeling packet losses as a Bernoulli process. They established statistical convergence properties of the expected covariance of estimation error and identified a critical observation-arrival rate, below which the covariance diverges and above which it remains bounded.
@@
Gupta et al. [gupta2006stochastic] extended this random measurement framework to a more general sensor scheduling problem. They proposed a stochastic sensor selection strategy that activates one sensor at each time step according to a prescribed probability distribution. Using the discrete-time Riccati recursion, they analyzed the evolution of the covariance of estimation error in expectation, deriving upper and lower bounds. Their results further demonstrated, through examples, that these bounds accurately approximate the actual steady-state covariance.
@@
Subsequent research generalized these results to the setting of random measurement matrices. For example, Marelli et al. [marelli2019stability] considered the case where both the measurement matrix and the noise covariance are drawn randomly at each time from a known distribution, and established necessary and sufficient conditions for mean-square stability of the KF. While closely related to our problem, their analysis is limited to stability criteria and does not provide bounds or steady-state characterizations of the error covariance.
@@
Calle and Bopardikar [calle2021probabilistic] derived probabilistic semidefinite bounds on the covariance of estimation error for randomized sensor selection in Kalman Filtering. Their framework, however, models sensor selection as a one-time random sampling of a subset of sensors that remains fixed thereafter, in contrast to the per-step stochastic scheduling analyzed by Gupta et al. [gupta2006stochastic].
@@
Early approaches primarily relied on the discrete-time Riccati recursion to track the distribution or expectation of the covariance of estimation error. More recently, efforts have been made to analyze continuous-time process models with discrete-time measurement models [ahdab2025optimal, dasgupta2025steady, tanwani2020error]. For example, Ahdab et al. [ahdab2025optimal] modeled measurement arrivals as independent Poisson processes with sensor-specific rates and derived an upper bound on the mean posterior covariance of the continuous–discrete KF along the mean auxiliary state. Similarly, Dasgupta and Tanwani [dasgupta2025steady] studied continuous-time linear systems with Poisson-sampled observations. They showed that the Kalman–Bucy filter yields piecewise-deterministic covariance trajectories, and derived conditions on the mean observation rate that ensure bounded and convergent expected covariance of estimation error. Their results extend the discrete-time “dropout threshold” in [sinopoli2004kalman]. However, their setting assumes a constant measurement matrix, with randomness arising from the irregular Poisson-distributed observation times rather than from the measurement matrix itself. In contrast, our focus is on a more general problem, akin to the framework of Marelli et al. [marelli2019stability], where the measurement matrix itself is random.
@@
Related work has also addressed continuous measurement models. For example, [le2010scheduling] study sensor scheduling for linear time-invariant systems. The authors derive a convex relaxation of the scheduling problem, yielding a performance lower bound. A lower bound reflects only the best-case performance and offers no guarantee on how the actual performance compares. In contrast, an upper bound constrains the worst-case covariance of the estimation error, providing a concrete performance guarantee. Additionally, [ha2019periodic] consider informative planning of sensing agents over an infinite time horizon for continuous-time linear systems. The covariance bound derived in this work tends to be conservative, as it omits the contribution of the measurement term.
@@
Analysis of the covariance of the estimation error naturally informs the design of the sensing system. Once stability conditions or performance bounds have been established, a practical question arises: how should sensors be deployed or activated to guarantee a prescribed level of estimation accuracy? This leads to sensor scheduling formulations, in which the active subset is selected over time. Prior work has addressed these problems through design optimization grounded in covariance (or its bounds), deriving policies that guide per–time-step sensor activation [zhao2014optimal, ahdab2025optimal]. Another typical problem is optimal sensor placement: given system parameters, one may wish to determine the minimum number of sensors or the required sensing quality needed to achieve a prescribed uncertainty threshold [tzoumas2016sensor]. Most prior work adopts covariance (or surrogates such as its log-determinant) as the performance metric and designs placement/scheduling policies accordingly. However, covariance measures are often difficult to interpret or compare across different sensing configurations, since they are unit-dependent and different scalarizations (e.g., trace, log-determinant) can rank designs inconsistently. To evaluate the estimation quality, we employ clarity, as introduced in [agrawal2023sensor, agrawal2024multi], which normalizes the differential entropy of a continuous random variable to the interval $[0,1]$. In this paper, we show that within our closed-form upper-bound framework, the fundamental limits of estimation performance can be explicitly characterized, enabling direct quantification of the sensing resources required to achieve a prescribed uncertainty threshold.
@@ III Problem Formulation
$\mathcal{R}_{G}(\bm{\Sigma})=\bm{\bm{A}}\bm{\Sigma}+\bm{\Sigma}\bm{\bm{A}}^{\top}+\bm{Q}_{c}-\bm{\Sigma}\bm{G}\bm{\Sigma},$ (6)
@@ IV-A Continuous-Time Upper Bound of Expected Covariance
The operator $\mathcal{R}_{G}$ is concave, that is, for any $\bm{\Sigma}_{1},\bm{\Sigma}_{2}\in\mathcal{S}_{++}^{n}$ and $\beta\in[0,1]$,
$\mathcal{R}_{G}(\beta\bm{\Sigma}_{1}\!+\!(1-\!\beta)\bm{\Sigma}_{2})\succeq\beta\mathcal{R}_{G}(\bm{\Sigma}_{1})\!+\!(1-\!\beta)\mathcal{R}_{G}(\bm{\Sigma}_{2})$ (9)
@@
$\mathcal{R}_{G}(\bm{\Sigma}_{1})-\mathcal{R}_{G}(\bm{\Sigma}_{2})=\tilde{\bm{A}}\bm{K}+\bm{K}\tilde{\bm{A}}^{\top}$ (10)
@@
$=\mathbb{E}[\mathcal{R}_{\bar{G}}(\bm{\Sigma}(t))]$ (16d)
$\preceq\mathcal{R}_{\bar{G}}(\mathbb{E}[\bm{\Sigma}(t)])$ (16e)
where Eq. (16a) holds by linearity of expectation; Eq. (16b) follows by substituting the expression for $\dot{\bm{\Sigma}}$; Eq. (16c) uses the assumed independence of $\bm{\Sigma}(t)$ and $\bm{G}(t)$; Eq. (16d) invokes the definition of the Riccati operator in Eq. (6), and Eq. (16e) follows from Jensen’s inequality together with the concavity of $\mathcal{R}_{G}$ established in Lemma 1.
@@
$\succeq\mathcal{R}_{\bar{G}}(\bm{\Delta}(t))-\mathcal{R}_{\bar{G}}(\bar{\bm{\Sigma}}(t))$
@@ IV-B Closed-form Solution of Upper Bound
$0\;=\;\bm{A}\bm{\Delta}_{\infty}+\bm{\Delta}_{\infty}\bm{A}^{\top}+\bm{Q}_{c}\bm{I}-\bm{\Delta}_{\infty}\bar{\bm{G}}\bm{\Delta}_{\infty},$ (22)
@@
Since $\bar{\bm{G}}\in\mathbb{S}_{++}^{n}$, it admits the eigen-decomposition $\bar{\bm{G}}=\bm{U}\bm{\Lambda}_{G}\bm{U}^{\top}$, where $\bm{U}$ is orthogonal and $\bm{\Lambda}_{G}$ is diagonal. Under the isotropic assumption $\bm{A}=a\bm{I}$ and $\bm{Q}_{c}=q_{c}\bm{I}$, Eq. (22) implies that $\bar{\bm{G}}$ and $\bm{\Delta}_{\infty}$ commute, i.e., $\bar{\bm{G}}\bm{\Delta}_{\infty}=\bm{\Delta}_{\infty}\bar{\bm{G}}$. By Theorem 4.5.15 (a) in [horn2012matrix], any two commuting symmetric matrices can be simultaneously diagonalized by the same orthogonal matrix. Hence, applying the same factorization to $\bm{\Delta}_{\infty}$, we have $\bm{\Delta}_{\infty}=\bm{U}\bm{\Lambda}_{\Delta}\bm{U}^{\top}$, where $\bm{\Lambda}_{\Delta}$ is diagonal. Therefore, the Eq. (22) becomes
@@
$=2a(\bm{U}\bm{\Lambda}_{\Delta}\bm{U}^{\top})+q_{c}\bm{I}-(\bm{U}\bm{\Lambda}_{\Delta}\bm{U}^{\top})(\bm{U}\bm{\Lambda}_{G}\bm{U}^{\top})(\bm{U}\bm{\Lambda}_{\Delta}\bm{U}^{\top})$
$=q_{c}\bm{I}+\bm{U}(2a\bm{\Lambda}_{\Delta}-\bm{\Lambda}_{\Delta}\bm{\Lambda}_{G}\bm{\Lambda}_{\Delta})\bm{U}^{\top}$
$=\bm{U}(q_{c}\bm{I}+2a\bm{\Lambda}_{\Delta}-\bm{\Lambda}_{\Delta}\bm{\Lambda}_{G}\bm{\Lambda}_{\Delta})\bm{U}^{\top}$
where the term inside the parentheses is a diagonal matrix. For the equation to hold, each diagonal entry must be zero. Therefore, let $\gamma_{i}=[\bm{\Lambda}_{\Delta}]_{ii}$, $\lambda_{i}=[\bm{\Lambda}_{G}]_{ii}$, and we have
@@
Thus, we have derived the eigenvalues of $\bm{\Delta}_{\infty}$. Using these eigenvalues, we have
$\bm{\Delta}_{\infty}$ $=\bm{U}\bm{\Lambda}_{\Delta}\bm{U}^{\top}$ (28)
$=\bm{U}\big(\operatorname{diag}(\gamma_{1},.,\gamma_{n})\big)\bm{U}^{\top}$
$=\bm{U}\big(a\bm{\Lambda}_{G}^{-1}+\bm{\Lambda}_{G}^{-1/2}(a^{2}\bm{I}+q_{c}\bm{\Lambda}_{G})^{1/2}\bm{\Lambda}_{G}^{-1/2}\big)\bm{U}^{\top}.$